# Ant Colony Sampling with GFlowNets for Combinatorial Optimization

### v1

@@ 6.4 Ablation Study
We also conducted an ablation study for four key training techniques introduced in Section 5: $\beta$ annealing, guided exploration, energy reshaping, and shared energy normalization. The ablation for $\beta$ annealing is conducted by fixing $\beta$ to the maximum value of the annealed $\beta$ schedule, while for other elements, we simply remove them from our algorithm. The empirical results for TSP200 and CVRP200 reported in Table 1 validate the effectiveness of our design choices.
Table 2: Experimental result on TSP. Results for methods with * are drawn from Ye et al. (2023), Cheng et al. (2023a), and Jin et al. (2023). The reported time is the average computation time for solving a single instance.
TSP200 TSP500 TSP1000
Method Obj. Gap(%) Time(s) Obj. Gap(%) Times(s) Obj. Gap(%) Time(s)
LKH-3 10.72 0.00 9.80 16.55 0.00 31.70 23.14 0.00 102
AM* 12.00 11.91 0.2 21.46 29.07 0.6 33.55 45.10 1.5
POMO* 10.89 1.62 0.10 20.57 24.40 0.6 32.90 42.30 4.1
POMO + EAS 10.75 0.30 120 18.25 10.27 300 30.77 32.96 600
POMO + SGBS 10.77 0.43 133 18.86 13.96 318 28.58 23.51 605
DIMES+MCTS* - - - 17.01 2.78 11 24.45 5.75 32
SO* 10.78 0.63 10 16.94 2.40 15 23.77 2.80 26
Pointerformer* 10.79 0.68 5.5 17.14 3.56 14 24.80 7.90 40
H-TSP* - - - - - - 24.65 6.62 0.3
ACO $(T=10)$ 10.91 1.76 2.1 17.38 5.02 14 24.76 6.99 57
DeepACO $(T=10)$ 10.77 0.45 2.5 16.84 1.75 15 23.78 2.77 66
GFACS $(T=10)$ 10.75 0.26 2.5 16.80 1.54 15 23.72 2.53 66
Table 3: Experimental result on CVRP. The reported time is the average computation time for solving a single instance.
CVRP200 CVRP500 CVRP1000
Method Obj. Gap(%) Time Obj. Gap(%) Time Obj. Gap(%) Time
LKH-3 28.02 0.00 2.4m 63.07 0.00 18.6m 119.52 0.00 2.1h
POMO 29.12 3.94 1s 68.69 8.90 2s 145.74 21.94 3s
POMO + EAS 27.99 -0.11 2.0m 64.76 2.67 5.0m 126.25 5.63 10.0m
POMO + SGBS 28.07 0.19 2.0m 65.44 3.75 5.2m 127.90 7.01 10.1m
SymNCO 29.15 4.03 1s 68.81 9.10 2s 141.82 18.66 3s
SymNCO + EAS 28.01 0.27 2.0m 64.63 2.47 5.0m 125.58 5.07 10.0m
SymNCO + SGBS 28.10 0.30 2.0m 65.40 3.69 5.2m 127.53 6.70 10.1m
ACO $(T=10)$ 28.45 1.55 13s 64.62 2.45 36s 122.82 2.76 1.3m
DeepACO $(T=10)$ 28.44 1.49 14s 64.49 2.25 36s 122.15 2.20 1.3m
GFACS $(T=10)$ 28.32 1.09 14s 64.19 1.76 36s 121.80 1.91 1.3m
In the domain of machine learning for combinatorial optimization, significant strides have been made using various approaches, including sampling-based methods (Sun et al., 2023), heatmap-guided searches (Sun & Yang, 2023; Qiu et al., 2022), and constructive methods that employ (Drakulic et al., 2023; Kool et al., 2018; Kwon et al., 2020). Our study compares our methods with RL-based approaches in TSP and CVRP for sizes $N=200,500,1000$, which do not require expert-labeled data. This comparison is particularly relevant as our unsupervised approach aligns with sequence decision-making processes akin to RL.
Baselines. For TSP baselines, we select the AM (Kool et al., 2018), POMO (Kwon et al., 2020), Sym-NCO (Kim et al., 2022b) and Pointerformer (Jin et al., 2023) as RL-based constructive methods, DIMES (Qiu et al., 2022) as RL-based heatmap-based methods, the Select and Optimize (SO) (Cheng et al., 2023b) as RL-based improvement methods, and H-TSP (Pan et al., 2023b) as a RL-based hierarchical methods. We also applied two search algorithms, the Efficient Active Search (EAS) (Hottung et al., 2021) and Simulation Guided Beam Search (SGBS) (Choo et al., 2022), to POMO for a more thorough comparison. For CVRP [^1] We also adopt TAM as a CVRP baseline; however, due to the inaccessibility of source code, we compare it using the reported result in the paper (Hou et al., 2022). See Section C.2.. We choose POMO and Sym-NCO as baselines with EAS and SGBS. Note that we restrict the duration of these searches to no more than 2, 5, and 10 minutes for problems with 200, 500, and 1000 nodes, respectively. We use LKH-3 (Helsgaun, 2017) as an oracle heuristic solver for TSP and CVRP to obtain near-optimal solutions that serve as baselines with $0.00\%$ gap. Please refer to Section A.4 for the parameter settings of LKH-3.
Results. Among the baseline methods evaluated, our GFACS algorithm consistently delivers either superior or highly competitive performances. When compared to faster methods like AM, POMO or H-TSP, GFACS demonstrates substantially better results. Even after an additional search with EAS or SGBS, they fail to match GFACS’s achievements, except in CVRP200. Compared to algorithms like DIMES and Pointformer, which operate at similar speeds, GFACS achieves improved performance. GFACS performs slightly worse than the oracle heuristic solver LKH-3, but offers an advantage of faster inference, especially in CVRP. Therefore, we can conclude that GFACS, as a general-purpose algorithm, holds significant appeal in the highly competitive benchmarks of both TSP and CVRP.
7 Conclusion
This paper proposes a novel neural-guided meta-heuristic algorithm called generative flow ant colony sampler (GFACS). Our algorithm was designed by integrating generative flow networks (GFlowNets) and ant colony optimization (ACO) with a novel combination of training techniques. GFACS consistently outperforms other ACO methods in several representative combinatorial optimization tasks. GFACS also shows competitive results both in traveling salesman problems and capacitated vehicle routing problems compared to a range of RL-based methods.
@@ A.1 Learning architecture and optimizer
We utilize the anisotropic graph neural network with edge gating mechanisms, which is proposed by (Joshi et al., 2021) and adopted by (Ye et al., 2023), as the foundational GNN architecture for the construction of a heuristic matrix. This network offers node and edge embeddings that are more beneficial for edge prediction than other popular GNN variants, such as Graph Attention Networks (Veličković et al., 2018) or Graph Convolutional Networks (Kipf & Welling, 2016).
@@
$h_{i}^{\ell+1}=h_{i}^{\ell}+\alpha(\text{BN}(U^{\ell}h_{i}^{\ell}+\mathcal{A}_{j\in\mathcal{N}_{i}}(\sigma(e^{\ell}_{ij})\odot V^{\ell}h_{j}^{\ell})))$ (13)
@@
$e_{ij}^{\ell+1}=e_{ij}^{\ell}+\alpha(\text{BN}(m^{\ell}))$ (15)
where $U^{\ell},V^{\ell},P^{\ell},Q^{\ell},R^{\ell}\in R^{d\times d}$ are the learnable parameters of layer $\ell$, $\alpha$ denotes the activation function, BN(Ioffe & Szegedy, 2015) denotes batch normalization, $\mathcal{A}$ denote the aggregation function, $\sigma$ is the sigmoid function, $\odot$ is Hadamard product, $\mathcal{N}_{i}$ denotes the neighborhoods of node $v_{i}$. In this paper we use $\alpha$ is SiLU(Elfwing et al., 2018) and $\mathcal{A}$ is mean pooling. We transform the extracted edge features into real-valued heuristic metrics by employing a 3-layer Multilayer Perceptron (MLP) that includes skipping connections to the embedded node feature $h$. The SiLU is the activation function for all layers except the output layer, in which the sigmoid function is applied to generate normalized outputs. The training process is optimized using the AdamW optimizer(Loshchilov & Hutter, 2017) in conjunction with a Cosine Annealing Scheduler(Loshchilov & Hutter, 2016)
@@ A.2 Hyperparameters
TSP200 TSP500 TSP1000 CVRP200 CVRP500 CVRP1000 OP100 PCTSP100 SOP100 SMTWTP100 BPP120
$N_{\text{epoch}}$ 50 50 20* 50 50 20* 20 20 20 20 20
$N_{\text{inst}}$ 400 400 200* 200 200 100* 100 100 100 100 100
$M$ 20 20 10 10 10 5 5 5 5 5 5
$K$ 30 30 20 20 20 15 30 30 30 30 30
There are other hyperparameters to be considered specifically for GFACS. $\beta$ is one of the most important hyperparameters for training GFACS, which is analyzed in Section 6.3. As we’ve discussed in Section 5.1, we use annealing for $\beta$ using the following log-shaped scheduling:
@@
Table 5: Settings of $\beta$ for each task.
TSP CVRP OP PCTSP SOP SMTWTP BPP
$\beta_{\text{min}}$ 200 500 5 20 10 10 500
$\beta_{\text{max}}$ 1000 2000 10 50 20 20 5000
@@ A.3 Computing resource
A.4 Hyperparameters for LKH-3.
LKH-3 (Helsgaun, 2017), an extension of the Lin-Kernighan heuristic (Lin & Kernighan, 1973), is a powerful open-source heuristic solver designed for solving routing problems. It not only addresses TSP but also handles a wide array of constrained routing problems, such as CVRP and PDP, by minimizing a penalty term that measures constraint violations.
In the following, we describe the parameter settings for LKH-3 used in this work. For both TSP and CVRP, we set MAX_TRIALS to be $100\times N$ where $N$ is the number of nodes and SEED to $1234$. We set RUNS to 10 for TSP but 1 for CVRP to limit the search time in a reasonable range. For other parameters not stated, we followed the default settings.
@@ Appendix B Problem Settings
In this section, we provide an overview of the combinatorial optimization problems we address. These encompass vehicle routing problems, bin packing problems, and scheduling problems. For a comprehensive understanding of each problem’s specific settings, we direct readers to the work by Ye et al. (2023). Furthermore, for access to the exact source code pertaining to each problem, please refer to the repository available at  [^2] https://github.com/henry-yeh/DeepACO from Ye et al. (2023).
@@
An OP instance comprises a set of nodes to visit, each with its position and prize, a depot, and a maximum length constraint. We sample a random OP instance by first sampling the position of nodes, including the starting node, from a unit square $[0,1]^{2}$. And we set the prize for each node, following (Fischetti et al., 1998; Ye et al., 2023), to be $p_{i}=(1+\lfloor 99\cdot\frac{d_{0,i}}{\max_{j=1}^{n}{d_{0,j}}}\rfloor)/100$, where $d_{0,i}$ is the Euclidean distance from the depot to node $i$. We set the maximum length to 4 in the OP100 task, following Kool et al. (2018).
@@
Sequential Ordering Problem
The Sequential Ordering Problem (SOP) seeks to find a linear order of jobs that minimizes total processing time with a set of precedence constraints among the jobs. An SOP instance is a set of jobs and precedence constraints. Each job has its own processing time, and each pair has a waiting (transition) time.
We sample an SOP instance by sampling the processing time and waiting time from $U[0,1]$. The precedence constraints for each job pair are laid with probability 0.2, while ensuring feasibility, i.e., when $i<j$ means that the constraint that job $i$ precedes the job $j$, the feasibility of a problem requires $i<j$ and $j<k$ then $i<k$.
@@
The Bin Packing Problem (BPP) involves packing objects of different volumes into a finite number of bins or containers of fixed capacity to minimize the number of bins used. In this work, we use the latter setting.
@@
Appendix C Additional Experiments
C.1 Experiments on TSPLIB
This section presents a comparison of performance between the baseline models and our model on real-world Traveling Salesman Problem (TSP) instances. We evaluated the models on TSP200, TSP500, and TSP1000 under their respective conditions. Our model outperforms the other Ant Colony Optimization (ACO) baselines, demonstrating superior generalization capabilities.
Table 6: Evaluation of ACO methods on the real-world TSP problem. ‘#’ denotes the number of instances in each set.
Size # ACO DeepACO Ours
$300100\leq N<$ 300 30 1.71% 1.25% 1.21%
$700300\leq N<$ 700 10 4.26% 2.70% 2.60%
$1500700\leq N<$ 1500 12 7.01% 4.08% 3.88%
ALL 52 3.43% 2.24% 2.16%
C.2 Comparison against TAM
We compare GFACS and DeepACO against CVRP instances used in TAM (Hou et al., 2022). These instances have a different capacity constraint from those we used for the main experiment. Please refer to (Hou et al., 2022) for more details regarding the instance generation. Moreover, to match the processing time to be comparative to DeepACO, we set $K$ (equal to the number of ants for ACO) to 16.

### v2

@@ 6.3 Ablation study
Table 3: Results of the ablation study for the proposed training techniques.
Method TSP500 CVRP500 BPP120
GFACS 16.80 64.19 51.84
   – Off-policy Training 17.09 64.42 52.32
   – Energy Reshaping 16.87 64.41 52.07
   – Shared $\varepsilon$-norm 17.52 65.51 90.73
    + VarGrad [82] 16.86 64.28 51.95
This section reports the ablation study results, validating our proposed training techniques for stabilizing and enhancing GFlowNets in large-scale combinatorial optimization tasks. As detailed in Table 3, our off-policy training method, incorporating local search, energy reshaping to value potential solutions, and shared $\epsilon$-norm to stabilize conditional GFlowNets training, significantly improves performance across three tasks: TSP, CVRP, and BPP. For comprehensive results, please refer to Table 10.
6.4 Comparative analysis with previous GFlowNets training methods
While this paper proposes a novel training method for GFlowNets designed for large-scale combinatorial optimization problems that are not locally decomposable due to their cyclic nature (e.g., TSP, CVRP) [1], we also compare existing GFlowNets training methods within this domain.
Figure 4: Validation costs during training with different GFlowNet losses.
Forward-looking detailed balance (FL-DB) [59, 84]. FL-DB leverages local credits when the application benefits from useful local rewards for partial solutions. While graph combinatorial optimization tasks are generally locally decomposable [1], yielding useful local rewards, tasks like TSP do not offer substantial local rewards due to their cyclic nature, which imposes a global constraint, thus complicating local credit assignment. As shown in Figure 4, various GFlowNet methods were used to train the ACO’s heuristic matrix $\eta(\mathcal{G};\theta)$. While FL-DB did not surpass on-policy TB for TSP, our off-policy method demonstrated significant improvement. This trend holds across different deep learning models and inference tasks beyond ACO (see Section D.1).
Gradient of log-parition variance (VarGrad) [65, 82]. The VarGrad loss, initially proposed by Richter et al. [65] for variational inference, was applied to conditional GFlowNets for solving scheduling problems by Zhang et al. [82]. This approach is crucial for combinatorial optimization tasks that require conditional inference on problem instances ($\mathcal{G}$). The key idea involves using a generative distribution and energy to compute the log-partition function through the consistency property of log-partitions over trajectories conditioned on identical instances ($\mathcal{G}$), thereby bypassing the explicit training of the (conditional) log-partition ($\log Z(\mathcal{G})$). Explicit training of the log-partition is challenging in conditional GFlowNets due to the varying energy landscapes across different instance conditions ($\mathcal{G}$). On the other hand, our method, which utilizes explicit training of the log-partition (see Equation 9), outperforms VarGrad in training conditional GFlowNets (see Table 3). This improvement is achieved because our Share $\epsilon$-norms enable stable and explicit training of the log-partition by normalizing energy values into sample means, resulting in an energy distribution across different instances that maintains a mean of zero. Note that accurately estimated the explicit log-partition value is known to stabilize the amortized inference process in GFlowNets [53].
7 Discussion
This paper proposes a novel probabilistic search method leveraging the off-policy training capabilities of GFlowNets. This strategy can be applied not only to ant-colony optimization (ACO) but also to other probabilistic methods requiring an effective learned prior (see Section D.4). Our off-policy framework enables the integration of learning algorithms with offline datasets containing optimal combinatorial solutions for specific problem instances, combined with online exploration to fine-tune the policy through off-to-on training. This approach holds promise for future work.
Limitation. A limitation of our method is that its performance depends on the effectiveness of the local search. A promising future direction could involve leveraging advanced neural-guided improvement heuristics, including neural local search approaches [48, 50], to create a hybrid algorithm that combines GFlowNets with neural improvement-based local search.
@@ A.1 Learning architecture and optimizer
We utilize the anisotropic graph neural network with edge gating mechanisms, which is proposed by [29] and adopted by [79], as the foundational GNN architecture for the construction of a heuristic matrix. This network offers node and edge embeddings that are more beneficial for edge prediction than other popular GNN variants, such as Graph Attention Networks [73] or Graph Convolutional Networks [40].
@@
$h_{i}^{\ell+1}=h_{i}^{\ell}+\text{ACT}(\text{BN}(U^{\ell}h_{i}^{\ell}+\mathcal{A}_{j\in\mathcal{N}_{i}}(\sigma(e^{\ell}_{ij})\odot V^{\ell}h_{j}^{\ell})))$ (12)
@@
$e_{ij}^{\ell+1}=e_{ij}^{\ell}+\text{ACT}(\text{BN}(m^{\ell}))$ (14)
where $U^{\ell},V^{\ell},P^{\ell},Q^{\ell},R^{\ell}\in R^{d\times d}$ are the learnable parameters of layer $\ell$, ACT denotes the activation function, BN[25] denotes batch normalization, $\mathcal{A}$ denote the aggregation function, $\sigma$ is the sigmoid function, $\odot$ is Hadamard product, $\mathcal{N}_{i}$ denotes the neighborhoods of node $v_{i}$. In this paper we use $\alpha$ is SiLU[17] and $\mathcal{A}$ is mean pooling. We transform the extracted edge features into real-valued heuristic metrics by employing a 3-layer Multilayer Perceptron (MLP) that includes skipping connections to the embedded node feature $h$. The SiLU is the activation function for all layers except the output layer, in which the sigmoid function is applied to generate normalized outputs. The training process is optimized using the AdamW optimizer[46] in conjunction with a Cosine Annealing Scheduler[45]
@@ A.2 Hyperparameters
TSP200 TSP500 TSP1000 CVRP200 CVRP500 CVRP1000 OP100 PCTSP100 SMTWTP100 BPP120
$N_{\text{epoch}}$ 50 50 20* 50 50 20* 20 20 20 20
$N_{\text{inst}}$ 400 400 200* 200 200 100* 100 100 100 100
$M$ 20 20 10 10 10 5 5 5 5 5
$K$ 30 30 20 20 20 15 30 30 30 30
@@
There are other hyperparameters to be considered specifically for GFACS. $\beta$ is one of the most important hyperparameters for training GFACS, which is analyzed in Section D.2. We use annealing for $\beta$ using the following log-shaped scheduling:
@@
Table 5: Settings of $\beta$ for each task.
TSP CVRP CVRPTW OP PCTSP SMTWTP BPP
$\beta_{\text{min}}$ 200 500 500 5 20 10 1000
$\beta_{\text{max}}$ 1000 2000 2000 10 50 20 2000
@@ Appendix B Problem settings
In this section, we provide an overview of the combinatorial optimization problems we address. These encompass vehicle routing problems, bin packing problems, and scheduling problems. For a comprehensive understanding of each problem’s specific settings, we direct readers to the work by Ye et al. [79]. Furthermore, for access to the exact source code pertaining to each problem, please refer to the repository available at  [^3] https://github.com/henry-yeh/DeepACO from Ye et al. [79].
@@
Capacitated vehicle routing problem with time window
A uniform-capacity fleet of delivery vehicles is tasked with servicing customers who have specific demand and operating hours (earliest-latest arrival) for a single commodity. The vehicles should deliver to customer only the time available for service. The goals are to minimize the total travel distance.
A CVRPTW instance consists of a set of customer nodes, a single depot node, the demands of each customer node, the vehicle capacity (as in CVRP), and additionally, the earliest and latest arrival times for each node, as well as the service time at each node. We generate the CVRPTW instance following the method provided by Berto et al. [6], which can be accessed here [^4] https://github.com/ai4co/rl4co/blob/main/rl4co/envs/routing/cvrptw/generator.py.
@@
An OP instance comprises a set of nodes to visit, each with its position and prize, a depot, and a maximum length constraint. We sample a random OP instance by first sampling the position of nodes, including the starting node, from a unit square $[0,1]^{2}$. And we set the prize for each node, following [19, 79], to be $p_{i}=(1+\lfloor 99\cdot\frac{d_{0,i}}{\max_{j=1}^{n}{d_{0,j}}}\rfloor)/100$, where $d_{0,i}$ is the Euclidean distance from the depot to node $i$. We set the maximum length to 4 in the OP100 task, following Kool et al. [41].
@@
The bin packing problem (BPP) aims to minimize the number of bins when objects with different volumes need to be packed into a finite number of bins. The cost functions are often modified to maximize the utility of used bins [18, 79]. However, the direct use of the objective function employed in DeepACO can lead to an excessively large $\beta$ since the objective values tend to be saturated near 1.0; note that the objective values can not exceed 1.0 in that setting. Therefore, we slightly modify them to minimize a weighted sum of the inverse of the utilization of the used bins throughout all our BPP experiments.
@@
Appendix C Implementation of destroy-and-repair local search
As demonstrated in Section 5, our algorithm utilizes the local search for off-policy training. The refined solutions from local search serve as exploitation batches, and they are also used for energy-reshaping (Eq. 7). For some problems like traveling salesman problem (TSP), capacitated vehicle routing problem (CVRP), and CVRP with time window (CVRPTW), we used the well-established local search operators such as 2-opt, Swap* [74] and granular neighborhood search from PyVRP [77].
While domain-specific local search techniques are well developed for many combinatorial optimization problems, there are still tasks that lack effective local search solvers like Swap*. A potential candidate for a general-purpose local search solver in combinatorial optimization is the destroy-and-reconstruct local search operator. This method is inspired by the large neighborhood search approach described by Pisinger and Ropke [61]. Our implementation utilizes an evolutionary-style algorithm that repeatedly performs destruction, reconstruction, and top-K selection over a specified number of rounds.
Symmetricity-aware destruction. The most naive way to destroy a solution $x$ is simply dropping $N_{destroy}$ of the last nodes from the sampled trajectory $\tau_{\to x}$. However, it only performs the destruction with one trajectory, while there can be multiple trajectories that correspond to the same solution $x$. Considering this symmetric nature, our destruction phase consists of 1) selecting $\tau^{\prime}_{\to x}$ randomly from the set of symmetric trajectories and then 2) destroying $\tau^{\prime}_{x}$ by dropping the last $N_{destroy}$ nodes, resulting in a partial solution. Note that the number of symmetric trajectories for a given solution differs from problem types, as discussed in Appendix E.
Reconstruction with temperature annealing. Given a partial solution, we reconstruct a new solution $x^{\prime}$ using the policy similar to the decoding steps explained in Section 4.2. Considering the purpose of the local search (refining a given solution), we further promote the exploitation by introducing an inverse temperature for the policy. Formally, we modify Equation 5 with the inverse temperature $\delta_{r}$ as follows:
$P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho,\delta_{r})\propto\begin{cases}\left(\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G};\theta)\right)^{\delta_{r}}&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}$ (15)
We increase $\delta_{r}$ linearly from 1.0 to a specified maximum value, $\delta_{\text{max}}$, as the local search round progresses.
Top-K selection. For a given solution $x$, we perform the aforementioned destruction and reconstruction multiple times (in batch), obtaining a set of refined solutions. Then we evaluate them and leave only top-K solutions among them in terms of solution quality. The remaining solutions are used as an initial solution for the next round of the search.
Appendix D Additional experiments
D.1 Comparison between trajectory balance and forward-looking detailed balance
We trained Attention Model (AM) [41] and heatmap-based policy from Joshi et al. [29] in an unsupervised way using FL-DB and TB loss. Table 6 shows the test results for 128 random TSP50 instances. The result is aligned with the previous observation from GFACS, showing the limitation of FL-DB loss for solving complex CO tasks.
Table 6: Comparison between forward-looking detailed balance (FL-DB) and trajectory balance (TB) loss in TSP50, using Attention Model (AM) and heatmap-based method from Joshi et al. [29] as backbone policies.
TSP50
Method Obj. Gap (%)
Concorde 5.69 0.00
[…]
D.2 Analysis of $\beta$ and diversity-optimality trade-off
In this subsection, we explored the diversity-optimality trade-off by varying the inverse energy temperature $\beta$ from low to high values. For analysis, we generate $K$ solution samples from policies trained with GFACS with different values of $\beta$ and DeepACO (with $\rho=1$ so that the pheromone does not affect the result).
We use average pairwise Jaccard distance as a diversity metric. Formally, the diversity of a set of $K$ solutions is calculated as $\frac{1}{K(K-1)}\sum_{i}\sum_{j}d_{J}(x_{i},x_{j})$, where $d_{J}$ is a Jaccard distance between two solutions, defined as follows:
$E(x)$ is the set of edges composing the solution $x$. In TSP or CVRP, the (undirected) edge is a set of adjoining nodes in the solution path. For some CO problems, such as SMTWTP, the edge should be directed due to its asymmetric nature.
Figure 5: The trade-off between diversity and optimality as the inverse energy temperature $\beta$ changes.
Figure 5 show that lower $\beta$ values lead to more diverse solutions with higher costs, while higher $\beta$ values result in less diverse solutions with lower costs. GFACS confirms the anticipated trade-off between diversity and optimality, establishing a Pareto frontier over DeepACO by offering more diverse and high-quality solutions. This advantage is particularly beneficial when an effective local search mechanism is available, increasing the likelihood of achieving superior local optima. However, excessively high $\beta$ (e.g., $2000$ in our result) degrades both diversity and solution quality, supposedly because too much high $\beta$ hinders exploration during training. This empirical evidence supports our design choice of $\beta$ annealing Section A.2, which encourages exploration in the early stage of training.
D.3 Evaluation with other ACO variants
Table 7: Evaluation results with different pheromone update rules, Elitist AS and MAX-MIN AS. We run $T=10$ ACO rounds with 100 ants in all settings.
TSP200 CVRP200
Method Obj. Gap (%) Obj. Gap (%)
Concorde / PyVRP 10.72 0.00 27.93 0.00
AS 10.91 1.77 28.45 1.86
DeepACO + AS 10.77 0.45 28.44 1.80
GFACS + AS 10.75 0.26 28.32 1.40
Elitist AS 10.97 2.31 28.46 1.90
DeepACO + Elitist AS 10.77 0.49 28.45 1.84
GFACS + Elitist AS 10.75 0.28 28.34 1.47
MAX-MIN AS 10.97 2.34 28.46 1.88
DeepACO + MAX-MIN AS 10.78 0.57 28.44 1.82
GFACS + MAX-MIN AS 10.75 0.32 28.33 1.43
Our method is focused on enhancing the design of the heuristic matrix within the ACO framework. Thus, the advancement of the ACO algorithm itself (usually in pheromone update and action selection rule) is orthogonal to our work. To validate this claim, we evaluate our method with different ACO variants other than AS, namely Elitist Ant System [14], MAX-MIN Ant System [68]. The results in Table 7 demonstrate that GFACS consistently outperforms both ACO with human-designed heuristic and DeepACO under all pheromone update rules tested. This underscores the superiority of GFACS in generating a high-quality heuristic matrix.
D.4 Applying our method to other search algorithms
Table 8: Evaluation results combine with active search as posterior inference method.
TSP200 CVRP200
Method Obj. Gap (%) Obj. Gap (%)
Concorde / PyVRP 10.72 0.00 27.93 0.00
Pretrained GFACS
   – Off-policy TB Active Search 10.75 0.29 28.34 1.47
   – On-policy RL Active Search 10.76 0.40 28.38 1.62
Pretrained DeepACO
   – On-policy RL Active Search 10.79 0.65 28.43 1.79
In this subsection, we study active search [3] rather than the pheromone update of ACO. In ACO, pheromone update is a posterior inference given a prior distribution (heuristic) for the probability model that generates the solutions. Similarly, we can think of active search as a posterior update that implicitly updates a given probability model by adjusting the parameters that generate it. We sample 20 solutions and take 5 gradient steps to update the parameter at each iteration. After 50 iterations, we report the average performance of the active search using our method with 3 seeds of 128 instances for each task. Table 8 shows that GAFCS integrates better with active search than DeepACO. This indicates that GAFCS is effective not only with heuristic-based posterior inference but also with learning-based posterior inference.
D.5 Comparison against TAM